OpenStudy (anonymous):

I'm taking Algrebra and struggling. Can someone PLEASE help me and show me how to do these two problems step by step. Question 1.) (x+2)^2=(x+y)^2=x^2+2xy+y^2 Question 2.) (x-3)^3=(x-y)^3=x^3-3x^2y+3xy^2-y^3

OpenStudy (sleepyjess):

let me know when you are back on and i will help you @cherrytree

OpenStudy (anonymous):

I'm back on. @sleepyjess

OpenStudy (sleepyjess):

im here

OpenStudy (sleepyjess):

ok so we have \((x+y)^2=x^2+2xy+y^2\)

OpenStudy (sleepyjess):

do you want to know how to get to that answer?

OpenStudy (anonymous):

yes please

OpenStudy (sleepyjess):

ok

OpenStudy (sleepyjess):

\((x+y)^2=(x+y)*(x+y)\) correct?

OpenStudy (anonymous):

yeah

OpenStudy (sleepyjess):

ok so we are going to use the FOIL method are you familiar with that?

OpenStudy (anonymous):

hmm somewhat

OpenStudy (sleepyjess):

 that is the F in FOIL multiply the first terms of each polynomial